# Obfuscation-Resilient Binary Code Similarity Analysis using Dominance Enhanced Semantic Graph

### v1

@@ 6.2. Resilience to Obfuscation in BCSA
Recent approaches have been addressing the challenges posed by obfuscation. Asm2vec (Ding et al., 2019) applies random walk and edge coverage on the CFG to sample instruction sequences, then trains PV-DM model to generate function embeddings. CoP (Luo et al., 2017) integrates program semantics with fuzzy matching of the longest common subsequence to capture basic block semantics. BinFinder (Qasem et al., 2023) first identifies a set of engineered interpretable features, then trains a neural network model using a siamese network architecture. GTrans (Zhang et al., [n. d.]) leverages the D-Tree structure and employs three encoding strategies to train a graph attention network. HermesSim (He et al., 2024) introduces a SOG to represent binary functions and generates function embeddings using GGNN. Our experimental results shows its resilience against obfuscation. Sem2vec (Wang et al., 2023) learns continuous and short execution traces from the CFG, extracts constraints through symbolic execution. Trex (Pei et al., 2023) learns program execution semantics from a mixture of regular and forced-execution traces, and then transfers the acquired semantic knowledge to match semantically similar binary functions. Although these approaches are resilient to obfuscation and can detect similarities between obfuscated binary, they still fall short of capturing the complete semantics of obfuscated binary, limiting their effectiveness when multiple obfuscation techniques are applied simultaneously.

### v2

@@ 6.2. Resilience to Obfuscation in BCSA
Recent approaches have been addressing the challenges posed by obfuscation. Asm2vec (asm2vec2019) applies random walk and edge coverage on the CFG to sample instruction sequences, then trains PV-DM model to generate function embeddings. CoP (semantics2017) integrates program semantics with fuzzy matching of the longest common subsequence to capture basic block semantics. BinFinder (binfinder2023) first identifies a set of engineered interpretable features, then trains a neural network model using a siamese network architecture. GTrans (gtrans2024) leverages the D-Tree structure and employs three encoding strategies to train a graph attention network. HermesSim (sog2024) introduces a SOG to represent binary functions and generates function embeddings using GGNN. Our experimental results shows its resilience against obfuscation. Trex (Trex2023) learns program execution semantics from a mixture of regular and forced-execution traces. Although these approaches are resilient to obfuscation and can detect similarities between obfuscated binary, they still fall short of capturing the complete semantics of obfuscated binary, limiting their effectiveness when multiple obfuscation techniques are applied simultaneously.
@@ 7. Conclusion
GenAI Disclosure Statement
No generative artificial intelligence (GenAI) tools are used in any stage of the research.
Acknowledgements.
The authors would like to appreciate the anonymous reviewers for their feedback. This project is supported by the Shenzhen Science and Technology Foundation (JCYJ20210324093212034), Guangdong Province Undergraduate University Quality Engineering Project (Shenzhen University Academic Affairs [2022] No. 7), National Natural Science Foundation of China (62272315), Science and Technology R&D Program of Shenzhen (20220810135520002), Guangdong Basic and Applied Basic Research Foundation (2023A1515011296), the Stable Support Project of Shenzhen (20231120145719001), Guangdong Province Key Laboratory of Popular High Performance Computers 2017B030314073, and Guangdong Province Engineering Center of China-made High Performance Data Computing System.